# The Energy Cascade Rate in Supersonic Magnetohydrodynamic Turbulence

### v1

@@ 2.1 Compressible MHD model
The governing equations for CMHD consist of the continuity equation for mass density, the momentum equation for the velocity field incorporating the Lorentz force, the induction equation for the magnetic field, and Gauss’s law for magnetism (see, e.g., Fitzpatrick, 2014; Andrés et al., 2017):
$\frac{\partial\rho}{\partial t}=-\boldsymbol{\nabla}\cdot(\rho\textbf{u}),$ (1)
$\frac{\partial\textbf{u}}{\partial t}=-\textbf{u}\cdot\boldsymbol{\nabla}\textbf{u}-\frac{\boldsymbol{\nabla}P}{\rho}+\frac{(\boldsymbol{\nabla}\times\textbf{B})\times\textbf{B}}{4\pi\rho}+\textbf{f}_{k}+\textbf{d}_{k},$ (2)
$\frac{\partial\textbf{B}}{\partial t}=\boldsymbol{\nabla}\times\left(\textbf{u}\times\textbf{B}\right)+\textbf{f}_{m}+\textbf{d}_{m},$ (3)
$\boldsymbol{\nabla}\cdot\textbf{B}=0,$ (4)
where $\rho$ represents the mass density, $\mathbf{u}$ is the velocity field, $P$ is the scalar pressure, and $\mathbf{B}$ denotes the total magnetic field. The equations include a mechanical forcing term and the curl of an electromotive forcing term ($\textbf{f}_{k}$ and $\textbf{f}_{m}$, respectively), alongside small-scale kinetic and magnetic dissipation terms ($\textbf{d}_{k}$ and $\textbf{d}_{m}$, respectively). The plasma is assumed to follow an isothermal equation of state, expressed as $P={c_{S}}^{2}\rho$, where $c_{S}$ is the constant sound speed. This relation allows the internal compressible energy of the plasma to be written as $e\equiv{c_{S}}^{2}\ln(\rho/\rho_{0})$, with $\rho_{0}$ being a reference density. The equations can be reformulated, as in Marsch & Mangeney (1987), in terms of internal energy by introducing the Alfvén velocity, $\mathbf{u}_{A}=\mathbf{B}/\sqrt{4\pi\rho}$, expressed in velocity units:
@@
$\textbf{u}_{\text{A}}\cdot\boldsymbol{\nabla}\rho=-2\rho(\boldsymbol{\nabla}\cdot\textbf{u}_{\text{A}}),$ (8)
where we have also defined the magnetic pressure, $P_{M}=\rho{u_{A}}^{2}/2$.
@@ 2.2.1 Flux terms
For incompressible MHD, the flux term can be recovered by neglecting density fluctuations (i.e., setting $\rho=\rho_{0}=1$). In this case, the flux term is given by:
$\textbf{F}_{I}=\langle\left[(\delta\mathbf{u})^{2}+(\delta\mathbf{u_{A}})^{2}\right]\delta\mathbf{u}-2(\delta\mathbf{u}\cdot\delta\mathbf{u_{A}})\delta\mathbf{u_{A}}\rangle$ (12)
. From these definitions of the flux terms it is possible to compute the compressible energy cascade rate, $\varepsilon_{F_{C}}$, and the incompressible cascade, $\varepsilon_{F_{I}}$ for each direction of the increments.
@@ 2.3 The energy cascade rate components
where $\text{S}_{j}$ corresponds to $\text{S}_{C}$, $\text{S}_{H}$, and $\text{M}_{\beta}$, respectively, along with their associated non-flux integrated terms. Finally, the total energy cascade rate can be expressed as the contribution from the flux term, $\varepsilon_{F_{C}}$, and the non-flux terms, $\varepsilon_{Q_{S}}$, $\varepsilon_{Q_{H}}$, and $\varepsilon_{Q_{M}}$. The total non-flux cascade is denoted by $\varepsilon_{Q}$.
@@ 3.1 Numerical code
In these simulations, first presented in Stalpes et al. (2024), turbulence is sustained by stochastic mechanical forcing designed to drive large-scale fluid motions. This forcing is initialized in Fourier space, where the acceleration field $\hat{\mathbf{f}}(\mathbf{k},t)$ evolves in real space according to a generalized Ornstein-Uhlenbeck process (Federrath et al., 2010). Energy is injected at a characteristic wavenumber corresponding to scales approximately half the domain size, ensuring the creation of large-scale eddies. The forcing combines both compressive and solenoidal components, with the input power distributed such that the ratio of compressive to solenoidal amplitude is 2/3. To emulate isothermal conditions, we set the adiabatic exponent to $\gamma=1.001$, ensuring the closure relation $P\simeq{c_{S}}^{2}\rho$. This setup allows the dissipation of kinetic energy to be redistributed into internal energy with minimal temperature variation, effectively approximating isothermal turbulence.
The forcing method ensures statistically isotropic turbulence. At large scales, the compressive component dominates, driving shock formation and small-scale instabilities. In contrast, the solenoidal component generates vorticity and sustains large-scale eddies. The driving field is helicity-free to machine precision, ensuring no parity violation.
@@ 3.2 Method
First, in Section 4.1, we present the parameter space and analyze field fluctuations as functions of the simulation inputs. In Section 4.2, we examine the signed energy cascade terms as a function of scale. Section 4.3 focuses on flux term analysis using the absolute increment method, which is widely employed in observational studies. Subsequently, we investigate the effects of compressibility and anisotropy in Sections 4.3.1 and 4.3.2, respectively. Finally, in Section 4.3.3, we provide a comprehensive comparison of the signed and absolute increment methods for both flux and non-flux contributions.
@@ 5 DISCUSSION AND CONCLUSIONS
We have performed three-dimensional direct numerical simulations to study the compressible and isothermal energy cascade rates in fully developed MHD turbulence across a broad range of plasma compressibility and magnetic guide field strengths. Specifically, we used the compressible and isothermal (Andrés et al., 2017; Banerjee & Galtier, 2013) exact relations to compute both the flux and non-flux contributions to the energy cascade rates. Our numerical results reveal a clear dependence of the energy cascade on plasma compressibility and the strength of the mean magnetic field. Moreover, we observed a consistent dominance of flux contributions over non-flux contributions when the signed cascade rate is computed. These findings highlight the role of compressibility and magnetic field anisotropy in modulating energy transfer mechanisms in compressible MHD turbulence.
The first estimation of the compressible energy cascade rate was obtained by Andrés et al. (2020) using PSP, THEMIS and MAVEN observations in the pristine solar wind. The authors computed both incompressible and compressible cascade rates for hundreds of events at different heliocentric distances, ranging from 0.2 au to 1.7, finding a moderate increases in $\varepsilon_{C}$ with respect to $\varepsilon_{I}$, depending on the compressibility and distance to the Sun. Our numerical results show that both compressible and incompressible flux terms increase as we increase the compressibility, with this effect being more pronounced as we increase the magnetic guide field strength. Furthermore, we observe that the purely compressible flux term, Eq. (11), is approximately an order of magnitude smaller than the Yaglom-like term, (10), in agreement with previous observation and numerical results (Banerjee et al., 2016; Hadid et al., 2017; Andrés et al., 2018, 2020; Brodiano et al., 2023). In particular, Brodiano et al. (2023) analyzed the signed and absolute values of incompressible and compressible MHD energy cascade rates in the solar wind at different heliocentric distances using PSP data and exact relations for compressible isothermal (Andrés et al., 2017) and polytropic (Simon & Sahraoui, 2022) MHD turbulence. Their observational results are fully consistent with our numerical results: a clear increase of the absolute value of $\langle|\varepsilon_{C}|\rangle$ and $\langle|\varepsilon_{I}|\rangle$ as the heliocentric distance decreases (i.e. as the magnetic guide field increases) and as the density fluctuations increase in the plasma.
@@
As we previously discussed, to calculate the energy cascade rates within the MHD scales using the exact law (9), we have employed two approaches: one based on signed increments of the variables and another using absolute increments. While the absolute value approach improves statistical convergence (e.g., Podesta et al., 2009), it tends to overestimate the non-flux contributions relative to the flux terms (see also, Brodiano et al., 2023; Andrés et al., 2019; Verdini et al., 2024). By contrast, the signed increment method, applied along different directions in the simulation domain, consistently highlights the dominance of the flux contributions across all levels of compressibility and mean magnetic field strengths explored in this work. However, it is important to note that this trend may not hold in localized regions of the plasma, such as current sheets or filamentary structures, where the hypothesis for the exact law derivation could not hold and presumably non-flux contributions could become dominant (Manzini et al., 2022; Ferrand et al., 2020). Specifically, Ferrand et al. (2020) have shown that in supersonic turbulence, a dominant compressible (non-flux) component can emerge, fundamentally differing from the conventional flux-driven cascade (Kolmogorov, 1941a; Politano & Pouquet, 1998a; Kritsuk et al., 2013) characterized by the development of filamentary structures. A detailed investigation of the compressible energy cascade rate in specific structures of the plasma in the MHD and ion scales is beyond the scope of this work.

### v2

@@ 2.1 Compressible MHD model
The governing equations for compressible magnetohydrodynamics (CMHD) include the continuity equation for mass density, the momentum equation for the velocity field with the Lorentz force, the induction equation for the magnetic field, and Gauss’s law for magnetism (see, e.g., Fitzpatrick, 2014; Andrés et al., 2017). Following Marsch & Mangeney (1987), the system can be expressed in terms of the velocity field $\mathbf{u}$, the Alfvén velocity $\mathbf{u}_{A}=\mathbf{B}/\sqrt{4\pi\rho}$ (where $\mathbf{B}$ is the magnetic field and $\rho$ the mass density), the scalar pressure $P$, and the internal energy $e$. Assuming an isothermal equation of state, the pressure is given by $P=c_{s}^{2}\rho$, where $c_{s}$ is the constant sound speed. Under this assumption, the internal compressible energy can be written as $e\equiv c_{s}^{2}\ln(\rho/\rho_{0})$, with $\rho_{0}$ a reference density:
@@
$\textbf{u}_{\text{A}}\cdot\boldsymbol{\nabla}\rho=-2\rho(\boldsymbol{\nabla}\cdot\textbf{u}_{\text{A}}).$ (4)
The model also includes a mechanical forcing term $\mathbf{f}_{k}$ and the curl of an electromotive forcing term $\mathbf{f}_{m}$, along with small-scale dissipation terms $\mathbf{d}_{k}$ and $\mathbf{d}_{m}$ acting on the velocity and magnetic fields, respectively. In addition, we define the magnetic pressure as $P_{M}=\rho u_{A}^{2}/2$.
@@ 2.2.1 Flux terms
For incompressible MHD, the flux term can be recovered by neglecting density fluctuations (i.e., setting $\rho=\rho_{0}=1$). In this case, the flux term is given by
$\textbf{F}_{I}=\langle\left[(\delta\mathbf{u})^{2}+(\delta\mathbf{u_{A}})^{2}\right]\delta\mathbf{u}-2(\delta\mathbf{u}\cdot\delta\mathbf{u_{A}})\delta\mathbf{u_{A}}\rangle.$ (8)
From these definitions of the flux terms it is possible to compute the compressible energy cascade rate, $\varepsilon_{F_{C}}$, and the incompressible cascade, $\varepsilon_{F_{I}}$ for each direction of the increments.
@@ 2.3 The energy cascade rate components
where $\text{S}_{j}$ corresponds to $\text{S}_{C}$, $\text{S}_{H}$, and $\text{M}_{\beta}$, respectively, along with their associated non-flux integrated terms. Finally, the total energy cascade rate can be expressed as the contribution from the flux term, $\varepsilon_{F_{C}}$, and the non-flux terms, $\varepsilon_{Q_{S}}$, $\varepsilon_{Q_{H}}$, and $\varepsilon_{Q_{M}}$. The total non-flux cascade, sum of these terms, is denoted by $\varepsilon_{Q}$.
@@ 3.1 Numerical code
In the numerical simulations, first presented in Stalpes et al. (2024), turbulence is reached and sustained by a stochastic mechanical forcing designed to drive large-scale fluid motions. This forcing is initialized in Fourier space, where the acceleration field $\hat{\mathbf{f}}(\mathbf{k},t)$ evolves in real space according to a generalized Ornstein-Uhlenbeck process (see also, Federrath et al., 2010). Energy is injected at a characteristic wavenumber corresponding to scales approximately half the domain size, ensuring the creation of large-scale eddies. Specifically, the forcing combines both compressive and solenoidal components, with the input power distributed such that the ratio of compressive to solenoidal amplitude is 2/3. To simulate isothermal conditions, we set the adiabatic exponent to $\gamma=1.001$, ensuring the closure relation $P\simeq{c_{S}}^{2}\rho$. This numerical setup allows the dissipation of kinetic energy to be redistributed into internal energy with minimal temperature variation, effectively approximating isothermal turbulence. The simulation boxes use periodic boundary conditions and a grid of $512^{3}$ points (or zones). The forcing method ensures statistically isotropic turbulence. At large scales, the compressive component dominates, driving shock formation and small-scale instabilities. In contrast, the solenoidal component generates vorticity and sustains large-scale eddies. The driving field is helicity-free to machine precision, ensuring no parity violation. Fully developed turbulence is reached after approximately five dynamical times, defined as $t_{dyn}=L_{0}/M_{S}$, where $L_{0}$ is the forcing scale (half the box size). This stage is identified by the saturation of the mean energy. More specifically, all analyses presented in this work are performed at $t=10t_{dyn}$ to ensure a clear stationary state.
All simulations exhibit field spectra with a Kolmogorov like $k^{-5/3}$ scaling and a well resolved inertial range, as shown in Stalpes et al. (2024). Notably, the slopes of the velocity and magnetic field spectra do not depend of $M_{S}$, whereas the density spectrum slope shows a strong positive correlation. The integral scale is roughly half the box, as expected, and the Taylor scale is approximately 20 computational zones. It is worth mentioning that in compressible turbulence the definition of the Taylor microscale remains valid under compressive driving because shear dissipation remains the dominant mechanism for which the Taylor scale was originally formulated. Although compressive effects introduce bulk viscous dissipation, the ratio of shear to compressive dissipation,
(×$\frac{\langle(\nabla\times\mathbf{v})^{2}\rangle}{\langle(\nabla\cdot\mathbf{v})^{2}\rangle}\sim 10^{4},$ (19)
is very large in our simulations. This indicates that shear effects overwhelmingly dominate energy dissipation, which justifies the use of the Taylor microscale as the relevant dissipation length scale.
@@ 3.2 Method
First, in Section 4.1, we present the parameter space and analyze field fluctuations as functions of the simulation inputs. In Section 4.2, we examine the signed energy cascade terms as a function of scale. Section 4.3 focuses on flux term analysis using the absolute increment method, which is widely employed in observational studies. Subsequently, we investigate the effects of compressibility and anisotropy in Sections 4.3.1 and 4.3.2, respectively. In Section 4.3.3 we provide a comprehensive comparison of the signed and absolute increment methods for both flux and non-flux contributions. Finally, in Section 4.3.4, we compare the estimated cascade to the energy injection rate.
@@ 4.3.2 Effect of anisotropy
4.3.4 Exact Relation vs. Energy Injection
Figure 7 (a) shows the energy injection rate, $\dot{E}$, as a function of the sonic Mach number, $M_{S}$, for all the runs. Figure 7 (b) shows the mean total compressible energy cascade rate, $\langle\varepsilon_{C}\rangle$, computed using the exact third-order law, plotted as a function of the energy injection rate, $\dot{E}$. In most cases, the cascade rate is found to be slightly lower than the injection rate. This discrepancy is likely due to the fact that $\varepsilon_{C}(\ell)$ is defined as a scale-dependent correlation function, and the estimate of $\langle\varepsilon_{C}\rangle$ relies on averaging over a finite portion of the inertial range. Such averaging may partially underestimate the true mean value. Despite this, a strong correlation is observed between the cascade and the injection rate.
Figure 7: (a) Energy injection rate, $\dot{E}$, as a function of the mean compressible energy cascade rate, $\langle\varepsilon_{C}\rangle$. (b) Mean energy cascade rate as a function of the energy injection rate. The color bar corresponds to the compressibility level, $\sigma_{\rho}$.
@@ 5 DISCUSSION AND CONCLUSIONS
We have performed three-dimensional direct numerical simulations to study the compressible and isothermal energy cascade rates in fully developed MHD turbulence across a broad range of plasma compressibility and magnetic guide field strengths. Specifically, we used the compressible and isothermal (Andrés et al., 2017; Banerjee & Galtier, 2013) exact relations to compute both the flux and non-flux contributions to the energy cascade rates. Our numerical results reveal a clear dependence of the energy cascade on plasma compressibility and the strength of the mean magnetic field. Moreover, we observed a consistent dominance of flux contributions over non-flux contributions when the signed cascade rate is computed in these compressibility levels.
An estimation of the compressible energy cascade rate was obtained by Andrés et al. (2020) using PSP, THEMIS and MAVEN observations in the pristine solar wind. The authors computed both incompressible and compressible cascade rates for hundreds of events at different heliocentric distances, ranging from 0.2 au to 1.7, finding a moderate increases in $\varepsilon_{C}$ with respect to $\varepsilon_{I}$, depending on the compressibility and distance to the Sun. Our numerical results show that both compressible and incompressible flux terms increase as we increase the compressibility, with this effect being more pronounced as we increase the magnetic guide field strength. Furthermore, we observe that the purely compressible flux term, Eq. (7), is approximately an order of magnitude smaller than the Yaglom-like term, (6), in agreement with previous observation and numerical results (Banerjee et al., 2016; Hadid et al., 2017; Andrés et al., 2018, 2020; Brodiano et al., 2023). In particular, Brodiano et al. (2023) analyzed the signed and absolute values of incompressible and compressible MHD energy cascade rates in the solar wind at different heliocentric distances using PSP data and exact relations for compressible isothermal (Andrés et al., 2017) and polytropic (Simon & Sahraoui, 2022) MHD turbulence. Their observational results are fully consistent with our numerical results: a clear increase of the absolute value of $\langle|\varepsilon_{C}|\rangle$ and $\langle|\varepsilon_{I}|\rangle$ as the heliocentric distance decreases (i.e. as the magnetic guide field increases) and as the density fluctuations increase in the plasma.
@@
The observed peak in the energy cascade rates at intermediate polar angles is consistent with prior findings in anisotropic turbulence studies such as Jiang et al. (2023). That study suggests that energy transfer rates vary significantly with polar angle due to differences in the effective inertial range across scales, as discussed by Wang et al. (2022). Our results support this notion, indicating that the cascade is most efficient at intermediate angles. One possible explanation is that, at these angles, energy transfer benefits from both strong perpendicular cascade dynamics (which dominate in quasi-2D turbulence) and contributions from parallel transfer mechanisms. This optimal alignment may maximize the rate at which energy is redistributed across scales before dissipation becomes dominant.
As we previously discussed, to calculate the energy cascade rates within the MHD scales using the exact law (5), we have employed two approaches: one based on signed increments of the variables and another using absolute increments. While the absolute value approach improves statistical convergence (e.g., Podesta et al., 2009), it tends to overestimate the non-flux contributions relative to the flux terms (see also, Brodiano et al., 2023; Andrés et al., 2019; Verdini et al., 2024). By contrast, the signed increment method, applied along different directions in the simulation domain, consistently highlights the dominance of the flux contributions across all levels of compressibility and mean magnetic field strengths explored in this work. However, it is important to note that this trend may not hold in localized regions of the plasma, such as current sheets or filamentary structures, where the hypothesis for the exact law derivation could not hold and presumably non-flux contributions could become dominant (Manzini et al., 2022; Ferrand et al., 2020). Specifically, Ferrand et al. (2020) have shown that in supersonic turbulence, a dominant compressible (non-flux) component can emerge, marking a fundamental departure from the traditional flux-driven cascade (Kolmogorov, 1941a; Politano & Pouquet, 1998a; Kritsuk et al., 2013), which is typically associated with filamentary structure formation. In this context, for sonic Mach numbers $M_{S}>6$, we expect the non-flux terms to become increasingly significant, as suggested by the trend observed in Figure 6, where the ratio of non-flux to flux contributions grows with increasing $M_{S}$. Figure 7 supports the cascade picture by showing correlation between the estimated total compressible energy cascade rate, $\langle\varepsilon_{C}\rangle$, and the energy injection rate. However, a detailed investigation of the compressible energy cascade rate within specific plasma structures, both at MHD and kinetic (ion) scales, remains an important direction for future work.